OpenStudy (anonymous):

The length of a table is 5 cm more than its width. Its area is 126 cm^2. Find its dimensions.

OpenStudy (anonymous):

L = 5+W A=W*L=126 W=? W ( W + 5 ) = 126 W^2 + 5W - 126 = 0 ( W - 9 ) ( W + 14 ) = 0 so W = 9 L=14